# An Accelerated Algorithm for Stochastic Bilevel Optimization under Unbounded Smoothness

### v1

@@ Appendix B Auxiliary Lemmas for Bilevel Optimization
By Lemma C.6, Lemma D.1 and $\mu$-strong convexity of $g$ in $y$, we have with probability at least $1-\delta$ over the randomness in ${\mathcal{F}}^{{\texttt{init}}}$ that
$\|y_{T_{0}}^{{\texttt{init}}}-y_{0}^{*}\|^{2}$ $\leq\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{T_{0}}U_{0}^{{\texttt{init}}}+\frac{\mu}{\mu+l_{g,1}}\frac{4\alpha\sigma_{g,1}^{2}}{\mu}\ln\frac{eT_{0}}{\delta}$
$\leq\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{T_{0}}U_{0}^{{\texttt{init}}}+\frac{\mu}{\mu+l_{g,1}}\frac{\epsilon^{2}}{16L_{0}^{2}}.$
@@
$U_{0}^{{\texttt{init}}}$ $=\frac{\mu}{2}\|y_{0}^{{\texttt{init}}}-y_{0}^{*}\|^{2}+g(x_{0},y_{0}^{{\texttt{init}}})-g(x_{0},y_{0}^{*})\leq\frac{\mu+l_{g,1}}{2}\|y_{0}^{{\texttt{init}}}-y_{0}^{*}\|^{2}.$
Now we set
$\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{T_{0}}\frac{\mu+l_{g,1}}{2}\|y_{0}^{{\texttt{init}}}-y_{0}^{*}\|^{2}+\frac{\mu}{\mu+l_{g,1}}\frac{\epsilon^{2}}{16L_{0}^{2}}\leq\frac{\mu}{\mu+l_{g,1}}\frac{\epsilon^{2}}{8L_{0}^{2}},$
@@
$T_{0}\geq\ln\left(\frac{\mu^{2}\epsilon^{2}}{8(\mu+l_{g,1})^{2}L_{0}^{2}\|y_{0}^{{\texttt{init}}}-y_{0}^{*}\|^{2}}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)$
By choice of $\alpha$ as in (35) and simple calculation we obtain $T_{0}=\widetilde{O}(1/\epsilon)$ when $\epsilon$ is small. ∎
@@
Under event ${\mathcal{E}}_{{\texttt{init}}}$, let $\{y_{t}\}$ be the iterates produced by Option I. Set $\alpha=\widetilde{\Theta}(\epsilon^{2})$ as in (35) and $\phi_{t}(y)=g(x_{t},y)=\frac{\mu}{2}\|y-y_{t}^{*}\|^{2}$ with $y\in\mathbb{R}$. Then for any $t\in[T]$, Algorithm 2 guarantees with probability at least $1-\delta$ over the randomness in ~$\widetilde{\mathcal{F}}_{T}^{1}$ (we denote this event as ${\mathcal{E}}_{y}^{1}$) that $\|y_{t}-y_{t}^{*}\|\leq\epsilon/2L_{0}$.
@@
$=\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{t}\left(\frac{\mu}{2}\|y_{T_{0}}^{{\texttt{init}}}-y_{0}^{*}\|^{2}+g(x_{0},y_{T_{0}}^{{\texttt{init}}})-g(x_{0},y_{0}^{*})\right)+\left(\frac{4\alpha\sigma_{g,1}^{2}}{\mu}+\frac{160\eta^{2}l_{g,1}^{2}}{\mu^{3}\alpha}\right)\ln\frac{eT}{\delta}$
$\leq\frac{\mu+l_{g,1}}{\mu}\|y_{T_{0}}^{{\texttt{init}}}-y_{0}^{*}\|^{2}+\frac{\epsilon^{2}}{16L_{0}^{2}}\leq\frac{3\epsilon^{2}}{16L_{0}^{2}}.$
@@
Under event ${\mathcal{E}}_{{\texttt{init}}}$, let $\{y_{t}\}$ be the iterates produced by Option II. set $\alpha^{\prime}=\mu\alpha/(\mu+l_{g,1})=\widetilde{\Theta}(\epsilon^{2})$ with $\alpha$ defined in (35) and run SNAG in each update round for
$N=\ln\left(\frac{1}{8}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)=\widetilde{O}(1/\epsilon)$ (38)
steps in every $I=\frac{\mu\epsilon}{2(1-\beta)L_{0}\sigma_{g,1}}=\widetilde{O}(1/\epsilon)$ iterations, set $\phi_{t}(y)=g(x_{t},y)$ when $t$ is a multiple of $I$ (i.e., $x_{t}$ is fixed for each update round of Option II so $g$ can be general functions). Then for any $t\in[T]$, Algorithm 2 guarantees with probability at least $1-\delta$ over the randomness in $\sigma(\cup_{t\leq T}\widetilde{{\mathcal{F}}}_{t}^{2})$ (we denote this event as ${\mathcal{E}}_{y}^{2}$) that $\|y_{t}-y_{t}^{*}\|\leq\epsilon/L_{0}$.
@@
At the beginning of the first round, by Lemma D.2 and D.3 we have $\|y_{0}-y_{0}^{*}\|\leq\epsilon/2L_{0}$, then we do not update the lower-level variable until $t=I$-th iteration, then for $t=I$, we have
@@
where we plug in the definition of $\eta$ and $I$. By $l_{g,1}$-smoothness of $g$ we have
$U_{I}$ $=\frac{\mu}{2}\|y_{I}-y_{0}^{*}\|^{2}+g(x_{I},y_{I})-g(x_{I},y_{I}^{*})\leq\frac{\mu+l_{g,1}}{2}\|y_{I}-y_{I}^{*}\|^{2}\leq\frac{\mu+l_{g,1}}{2}\frac{\epsilon^{2}}{L_{0}^{2}}$
@@
$\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{N}\frac{\mu+l_{g,1}}{2}\frac{\epsilon^{2}}{L_{0}^{2}}+\frac{\mu}{\mu+l_{g,1}}\frac{\epsilon^{2}}{16L_{0}^{2}}\leq\frac{\mu}{\mu+l_{g,1}}\frac{\epsilon^{2}}{8L_{0}^{2}},$
@@
$N\geq\ln\left(\frac{1}{8}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)$
By choice of $\alpha$ as in (35) and simple calculation we obtain $N=\widetilde{O}(1/\epsilon)$ when $\epsilon$ is small. Now we have
$\|y_{I+1}-y_{I}^{*}\|^{2}=\|y_{I}^{N}-y_{I}^{*}\|^{2}\leq\frac{\mu+l_{g,1}}{\mu}\|y_{I}-y_{I}^{*}\|^{2}+\frac{\epsilon^{2}}{16L_{0}^{2}}\leq\frac{3\epsilon^{2}}{16L_{0}^{2}},$
@@
$\|y_{I+1}-y_{I+1}^{*}\|\leq\|y_{I+1}-y_{I}^{*}\|+\|y_{I}^{*}-y_{I+1}^{*}\|\leq\frac{\sqrt{3}\epsilon}{4L_{0}}+\frac{\eta l_{g,1}}{\mu}\leq\frac{\epsilon}{2L_{0}},$
@@
Under 3.13.23.33.4 and event ${\mathcal{E}}_{{\texttt{init}}}\cap{\mathcal{E}}_{y}^{1}$ (Option I) or ${\mathcal{E}}_{{\texttt{init}}}\cap{\mathcal{E}}_{y}^{2}$ (Option II), we further set $\tau=\sqrt{\mu\alpha}$ in the averaging step (line 21 of Algorithm 2). Then for any $t\geq 0$ we have
@@
$\|\hat{y}_{0}-y_{0}^{*}\|=\|y_{T_{0}}^{{\texttt{init}}}-y_{0}^{*}\|\leq\sqrt{\frac{\mu}{8(\mu+l_{g,1})}}\frac{\epsilon}{L_{0}}\leq\frac{\epsilon}{L_{0}},$
@@
Under 3.13.23.33.4 and event ${\mathcal{E}}_{{\texttt{init}}}\cap{\mathcal{E}}_{y}^{1}$ (Option I) or ${\mathcal{E}}_{{\texttt{init}}}\cap{\mathcal{E}}_{y}^{2}$ (Option II), define $\epsilon_{t}=m_{t}-\mathbb{E}_{t}[{\bar{\nabla}f}(x_{t},\hat{y}_{t};\bar{\xi}_{t})]$, then we have the following averaged cumulative error bound:
@@
We apply Lemma B.4 and follow the similar procedure as in [34, Lemma D.9] to obtain
@@
where (46) follows by law of total expectation, (47) follows by Jensen’s inequality, (48) uses the fact that $\hat{\epsilon}_{j}$ for $j<t-i$ are ${\mathcal{F}}_{t-i-1}$-measurable, and are uncorrelated with $\hat{\epsilon}_{t-i}$; for (49) we use Lemma B.6 and D.5 to derive
@@
Suppose that 3.13.23.3 and 3.4 hold. For any $\eta$ satisfying
@@
where we use definition of $\epsilon_{t}$, Lemma B.4 and B.5 in the last inequality. Plugging the above inequality back into (53) we obtain
@@
Suppose 3.13.23.3 and 3.4 hold. Let $\{x_{t}\}$ be the iterates produced by Algorithm 2. For any given $\delta\in(0,1)$ and any small $\epsilon>0$ satisfying
@@
if we set parameters $\alpha,\alpha^{\prime},\alpha^{{\texttt{init}}},\beta,\gamma,\eta,\tau,I,N,S,Q,T_{0}$ as
@@
$\alpha^{{\texttt{init}}}=\frac{1-\beta}{\mu+l_{g,1}}\quad\alpha=\frac{1-\beta}{\mu},\quad\alpha^{\prime}=\frac{1-\beta}{\mu+l_{g,1}},\quad\gamma=\frac{1-\sqrt{\mu\alpha}}{1+\sqrt{\mu\alpha}},\quad\tau=1-\sqrt{\mu\alpha},$ (56)
$T_{0}=\ln\left(\frac{\mu^{2}\epsilon^{2}}{8(\mu+l_{g,1})^{2}L_{0}^{2}\|y_{0}^{{\texttt{init}}}-y_{0}^{*}\|^{2}}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right),$ (57)
$I=\frac{\mu\epsilon}{2(1-\beta)L_{0}\sigma_{g,1}},\quad N=\ln\left(\frac{1}{8}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right),$ (58)
@@
where $d_{0}$ and $P$ are defined in (36). Then with probability at least $1-2\delta$ over the randomness in $\sigma({\mathcal{F}}^{{\texttt{init}}}\cup\widetilde{{\mathcal{F}}}_{T}^{1})$ (for Option I) or $\sigma({\mathcal{F}}^{{\texttt{init}}}\cup(\cup_{t\leq T}\widetilde{{\mathcal{F}}}_{t}^{2}))$ (for Option II), Algorithm 2 guarantees $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|{\nabla\Phi}(x_{t})\|\leq 20\epsilon$ within $T=\frac{4d_{0}}{\eta\epsilon}=\widetilde{O}(1/\epsilon^{3})$ iterations, where the expectation is taken over the randomness in ${\mathcal{F}}_{T}$. For Option I, it is equivalent to $T_{0}+SQT=\widetilde{O}(1/\epsilon^{3})$ oracle calls of stochastic gradient or Hessian/Jacobian vector product. For Option II, it requires $T_{0}+\frac{NT}{I}+SQT=\widetilde{O}(1/\epsilon^{3})$ oracle calls.

### v2

@@ Appendix B Auxiliary Lemmas for Bilevel Optimization
By Lemmas C.6 and D.1 and $\mu$-strong convexity of $g$ in $y$, we have with probability at least $1-\delta$ over the randomness in ${\mathcal{F}}^{{\mathrm{init}}}$ that
$\|y_{T_{0}}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}$ $\leq\frac{2}{\mu}\left(1-\frac{\sqrt{\mu^{2}\alpha^{2}}}{4}\right)^{T_{0}}U_{0}^{{\mathrm{init}}}+\frac{4\mu\alpha^{2}\sigma_{g,1}^{2}}{\mu}\ln\frac{eT_{0}}{\delta}$
$\leq\frac{2}{\mu}\left(1-\frac{\mu\alpha}{4}\right)^{T_{0}}U_{0}^{{\mathrm{init}}}+\frac{\mu\alpha\epsilon^{2}}{64L_{0}^{2}}.$
@@
$U_{0}^{{\mathrm{init}}}$ $\leq\frac{2-\sqrt{\mu^{2}\alpha^{2}}+\mu^{2}\alpha^{2}}{2\mu\alpha^{2}}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}+g(x_{0},y_{0}^{{\mathrm{init}}})-g(x_{0},y_{0}^{*})$
$\leq\frac{3}{2\mu\alpha^{2}}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}+\frac{l_{g,1}}{2}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}\leq\frac{2}{\mu\alpha^{2}}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2},$
where the last inequality uses $\alpha\leq 1/l_{g,1}$. Now we set
$\frac{2}{\mu}\left(1-\frac{\mu\alpha}{4}\right)^{T_{0}}\frac{2}{\mu\alpha^{2}}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}+\frac{\mu\alpha\epsilon^{2}}{64L_{0}^{2}}\leq\frac{\mu\alpha\epsilon^{2}}{32L_{0}^{2}},$
@@
$T_{0}\geq\ln\left(\frac{\mu^{3}\alpha^{3}\epsilon^{2}}{256L_{0}^{2}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}}\right)\Big{/}\ln\left(1-\frac{\mu\alpha}{4}\right).$
By choice of $\alpha$ as in (35) and simple calculation we obtain $T_{0}=\widetilde{O}(\epsilon^{-2})$ when $\epsilon$ is small. ∎
@@
Under event ${\mathcal{E}}_{{\mathrm{init}}}$, let $\{y_{t}\}$ be the iterates produced by Option I. Set $\alpha=\widetilde{\Theta}(\epsilon^{2})$ as in (35) and $\phi_{t}(y)=g(x_{t},y)=\frac{\mu}{2}\|y-y_{t}^{*}\|^{2}$ with $y\in\mathbb{R}$. Then for any $t\in[T]$, Algorithm 2 guarantees with probability at least $1-\delta$ over the randomness in ~$\widetilde{\mathcal{F}}_{T}^{1}$ (we denote this event as ${\mathcal{E}}_{y}^{1}$) that $\|y_{t}-y_{t}^{*}\|\leq\epsilon/2L_{0}$.
@@
$\leq\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{t}\frac{2}{\alpha}\|y_{T_{0}}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}+\left(\frac{4\alpha\sigma_{g,1}^{2}}{\mu}+\frac{160\eta^{2}l_{g,1}^{2}}{\mu^{3}\alpha}\right)\ln\frac{eT}{\delta}$
$\leq\frac{4}{\mu\alpha}\|y_{T_{0}}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}+\frac{\epsilon^{2}}{64L_{0}^{2}}\leq\frac{\epsilon^{2}}{4L_{0}^{2}}.$
@@
Under event ${\mathcal{E}}_{{\mathrm{init}}}$, let $\{y_{t}\}$ be the iterates produced by Option II. Set $\alpha=\widetilde{\Theta}(\epsilon^{2})$ as in (35) and run SNAG in each update round for
$N=\ln\left(\frac{\mu\alpha}{128}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)=\widetilde{O}(\epsilon^{-1})$
steps in every $I=\frac{\mu\epsilon}{2(1-\beta)L_{0}\sigma_{g,1}}=\widetilde{O}(\epsilon^{-1})$ iterations, set $\phi_{t}(y)=g(x_{t},y)$ when $t$ is a multiple of $I$ (i.e., $x_{t}$ is fixed for each update round of Option II so $g$ can be general functions). Then for any $t\in[T]$, Algorithm 2 guarantees with probability at least $1-\delta$ over the randomness in $\sigma(\cup_{t\leq T}\widetilde{{\mathcal{F}}}_{t}^{2})$ (we denote this event as ${\mathcal{E}}_{y}^{2}$) that $\|y_{t}-y_{t}^{*}\|\leq\epsilon/L_{0}$.
@@
At the beginning of the first round, by Lemmas D.2 and D.3 we have $\|y_{0}-y_{0}^{*}\|\leq\epsilon/2L_{0}$, then we do not update the lower-level variable until $t=I$-th iteration, then for $t=I$, we have
@@
where in the last equality we plug in the definition of $\eta$ and $I$. By $l_{g,1}$-smoothness of $g$ we have
$U_{I}$ $=\frac{2-2\sqrt{\mu\alpha}+\mu\alpha}{2\alpha}\|y_{I}-y_{0}^{*}\|^{2}+g(x_{I},y_{I})-g(x_{I},y_{I}^{*})$
$\leq\frac{3}{2\alpha}\|y_{I}-y_{I}^{*}\|^{2}+\frac{l_{g,1}}{2}\|y_{I}-y_{I}^{*}\|^{2}\leq\frac{2\epsilon^{2}}{\alpha L_{0}^{2}}$
@@
$\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{N}\frac{2\epsilon^{2}}{\alpha L_{0}^{2}}+\frac{\epsilon^{2}}{64L_{0}^{2}}\leq\frac{\epsilon^{2}}{16L_{0}^{2}},$
@@
$N\geq\ln\left(\frac{\mu\alpha}{128}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)$
By choice of $\alpha$ as in (35) and simple calculation we obtain $N=\widetilde{O}(\epsilon^{-1})$ when $\epsilon$ is small. Now we have
$\|y_{I+1}-y_{I}^{*}\|^{2}=\|y_{I}^{N}-y_{I}^{*}\|^{2}\leq\frac{2}{\mu}\left(1-\frac{\sqrt{\mu\alpha}}{4}\right)^{N}\frac{2\epsilon^{2}}{\alpha L_{0}^{2}}+\frac{\epsilon^{2}}{64L_{0}^{2}}\leq\frac{\epsilon^{2}}{16L_{0}^{2}},$
@@
$\|y_{I+1}-y_{I+1}^{*}\|\leq\|y_{I+1}-y_{I}^{*}\|+\|y_{I}^{*}-y_{I+1}^{*}\|\leq\frac{\epsilon}{4L_{0}}+\frac{\eta l_{g,1}}{\mu}\leq\frac{\epsilon}{2L_{0}},$
@@
Under Assumptions 3.1, 3.2, 3.3 and 3.4 and event ${\mathcal{E}}_{{\mathrm{init}}}\cap{\mathcal{E}}_{y}^{1}$ (Option I) or ${\mathcal{E}}_{{\mathrm{init}}}\cap{\mathcal{E}}_{y}^{2}$ (Option II), we further set $\tau=\sqrt{\mu\alpha}$ in the averaging step (line 21 of Algorithm 2). Then for any $t\geq 0$ we have
@@
$\|\hat{y}_{0}-y_{0}^{*}\|=\|y_{T_{0}}^{{\mathrm{init}}}-y_{0}^{*}\|\leq\sqrt{\frac{\mu\alpha}{32}}\frac{\epsilon}{L_{0}}=\sqrt{\frac{1-\beta}{32}}\frac{\epsilon}{L_{0}}\leq\frac{\epsilon}{L_{0}},$
@@
Under Assumptions 3.1, 3.2, 3.3 and 3.4 and event ${\mathcal{E}}_{{\mathrm{init}}}\cap{\mathcal{E}}_{y}^{1}$ (Option I) or ${\mathcal{E}}_{{\mathrm{init}}}\cap{\mathcal{E}}_{y}^{2}$ (Option II), define $\epsilon_{t}=m_{t}-\mathbb{E}_{t}[{\bar{\nabla}f}(x_{t},\hat{y}_{t};\bar{\xi}_{t})]$, then we have the following averaged cumulative error bound:
@@
We apply Lemma B.4 and follow the similar procedure as in [35, Lemma D.9] to obtain
@@
where (45) follows by law of total expectation, (46) follows by Jensen’s inequality, (47) uses the fact that $\hat{\epsilon}_{j}$ for $j<t-i$ are ${\mathcal{F}}_{t-i-1}$-measurable, and are uncorrelated with $\hat{\epsilon}_{t-i}$; for (48) we use Lemmas B.6 and D.5 to derive
@@
Suppose that Assumptions 3.1, 3.2, 3.3 and 3.4 hold. For any $\eta$ satisfying
@@
where we use definition of $\epsilon_{t}$, Lemmas B.4 and B.5 in the last inequality. Plugging the above inequality back into (52) we obtain
@@
Suppose Assumptions 3.1, 3.2, 3.3 and 3.4 hold. Let $\{x_{t}\}$ be the iterates produced by Algorithm 2. For any given $\delta\in(0,1)$ and any small $\epsilon>0$ satisfying
@@
if we set parameters $\alpha,\alpha^{\prime},\alpha^{{\mathrm{init}}},\beta,\gamma,\eta,\tau,I,N,S,Q,T_{0}$ as
@@
$\alpha^{{\mathrm{init}}}=\frac{1-\beta}{\mu+l_{g,1}}\quad\alpha=\frac{1-\beta}{\mu},\quad\gamma=\frac{1-\sqrt{\mu\alpha}}{1+\sqrt{\mu\alpha}},\quad\tau=1-\sqrt{\mu\alpha},$ (55)
$T_{0}=\ln\left(\frac{\mu^{3}\alpha^{3}\epsilon^{2}}{256L_{0}^{2}\|y_{0}^{{\mathrm{init}}}-y_{0}^{*}\|^{2}}\right)\Big{/}\ln\left(1-\frac{\mu\alpha}{4}\right),$ (56)
$I=\frac{\mu\epsilon}{2(1-\beta)L_{0}\sigma_{g,1}},\quad N=\ln\left(\frac{\mu\alpha}{128}\right)\Big{/}\ln\left(1-\frac{\sqrt{\mu\alpha}}{4}\right),$ (57)
@@
where $d_{0}$ and $P$ are defined in (36). Then with probability at least $1-2\delta$ over the randomness in $\sigma({\mathcal{F}}^{{\mathrm{init}}}\cup\widetilde{{\mathcal{F}}}_{T}^{1})$ (for Option I) or $\sigma({\mathcal{F}}^{{\mathrm{init}}}\cup(\cup_{t\leq T}\widetilde{{\mathcal{F}}}_{t}^{2}))$ (for Option II), Algorithm 2 guarantees $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|{\nabla\Phi}(x_{t})\|\leq 20\epsilon$ within $T=\frac{4d_{0}}{\eta\epsilon}=\widetilde{O}(1/\epsilon^{3})$ iterations, where the expectation is taken over the randomness in ${\mathcal{F}}_{T}$. For Option I, it requires $T_{0}+SQT=\widetilde{O}(1/\epsilon^{3})$ oracle calls of stochastic gradient or Hessian/Jacobian vector product. For Option II, it requires $T_{0}+\frac{NT}{I}+SQT=\widetilde{O}(1/\epsilon^{3})$ oracle calls of stochastic gradient or Hessian/Jacobian vector product.